OpenStudy (anonymous):

Could somebody help me work through this? I totally forgot the steps! (2m^4 n^2/6m^2)^3

OpenStudy (anonymous):

\[(\frac{ 2m^4 n^2 }{ 6m^2 })^3\] A better representation of it

OpenStudy (anonymous):

@dan815 @bahrom7893 @jim_thompson5910 @primeralph would either of you guys be willing to help?

OpenStudy (anonymous):

@wio could you help? I know that first I need to solve within the parenthesis

OpenStudy (anonymous):

Cancel out the \(m\)s

OpenStudy (anonymous):

Wait do I cancel out the Ms first? because I simplified the 2/6 part first

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

\[ \frac{m^4}{m^2}=\frac{m^{\cancel 42}}{m^{\cancel 2 0}}=\frac{m^2}{1} \]

OpenStudy (anonymous):

I know how to cancel out m's...Do you see the big 2 and 6 there? I treated that as a seperate fraction sort of and simplified it first, was that wrong? Or should I have cancelled out the M's first? @wio

OpenStudy (anonymous):

It doesn't matter, mutliplication is the same order.

OpenStudy (anonymous):

Okay @wio so since the M's are cancelled out \[(\frac{ 2m^2n^2 }{ 6(1) })^3\]

OpenStudy (anonymous):

is that correct?@wio

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

yeah that is fine

OpenStudy (anonymous):

You can distribute the 3 at any time you want.

OpenStudy (anonymous):

Okay @wio \[\frac{ 2^3m^6n^6 }{ 6^3 }= \frac{ 8m^6n^6 }{ 216 }\]

OpenStudy (anonymous):

The 8 and the 216 can be reduced

OpenStudy (anonymous):

\[\frac{ m^6n^6 }{ 27 }\]

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

You can't simplify it anymore

OpenStudy (anonymous):

Okay @wio thank you!